OpenStudy (anonymous):

In the lab, Susan has two solutions that contain alcohol and is mixing them with each other. Solution A is 12% alcohol and Solution B is 40% alcohol. She uses 200 milliliters of Solution A. How many milliliters of Solution B does she use, if the resulting mixture is a 32% alcohol solution?

OpenStudy (anonymous):

Off the the top of my head... (% A)(mL A) + (% B)(mL B) = (% mixture)(mL mixture) (12% solute)(200mL) + (40% solute)(v mL) = (32% solute)((200+v) mL) You're solving for v. Anybody free to verify this answer? ^_^

OpenStudy (shane_b):

\[0.12(200)+0.40x=0.32(200+x)\]\[24+0.40x=64+0.32x\]\[0.40x-0.32x=64-24=40\]\[0.08x=40\]\[x=500\]So it will take 500mL of solution B to obtain a mixture of 32%.